OpenStudy (anonymous):

log 6x - log 3 = 4 I realize that logarithms are supposed to be easy...I honestly do not understand them. Laugh if you want

OpenStudy (anonymous):

they turn multiplication to addition\[\log(ab)=\log(a)+\log(b)\] and division to subtraction\[\log( \frac{ a }{ b })=\log(a)-\log(b)\]